# Bootstraping Clustering of Gaussians for View-consistent 3D Scene Understanding

### v1

@@ Abstract
Injecting semantics into 3D Gaussian Splatting (3DGS) has recently garnered significant attention. While current approaches typically distill 3D semantic features from 2D foundational models (e.g., CLIP and SAM) to facilitate novel view segmentation and semantic understanding, their heavy reliance on 2D supervision can undermine cross-view semantic consistency and necessitate complex data preparation processes, therefore hindering view-consistent scene understanding. In this work, we present FreeGS, an unsupervised semantic-embedded 3DGS framework that achieves view-consistent 3D scene understanding without the need for 2D labels. Instead of directly learning semantic features, we introduce the IDentity-coupled Semantic Field (IDSF) into 3DGS, which captures both semantic representations and view-consistent instance indices for each Gaussian. We optimize IDSF with a two-step alternating strategy: semantics help to extract coherent instances in 3D space, while the resulting instances regularize the injection of stable semantics from 2D space. Additionally, we adopt a 2D-3D joint contrastive loss to enhance the complementarity between view-consistent 3D geometry and rich semantics during the bootstrapping process, enabling FreeGS to uniformly perform tasks such as novel-view semantic segmentation, object selection, and 3D object detection. Extensive experiments on LERF-Mask, 3D-OVS, and ScanNet datasets demonstrate that FreeGS performs comparably to state-of-the-art methods while avoiding the complex data preprocessing workload. Our code is available at: https://github.com/wb014/FreeGS.
@@ Introduction
Figure 1: We present a novel framework, FreeGS, to inject 2D semantics (Refer to caption) into 3DGS (Refer to caption), without the need of any 2D labels. After a 2D-3D collaborative bootstrapping learning, the model can support versatile applications as shown.
@@
To resolve these challenges, we present FreeGS, an unsupervised 3DGS framework for view-consistent 3D scene understanding. The core of FreeGS is the IDentity-coupled Semantic Field (IDSF), which holds both semantic embeddings and view-consistent instance indices for each Gaussian. With only multi-view images as input, we optimize IDSF through a bootstrapping process that alternates between the Union-space 3D Gaussian Clustering phase and the Multi-level 2D Semantic Distillation phase. In the 3D clustering step, semantic information collaborates with the intrinsic properties of Gaussians (e.g., geometry and appearance) to extract unsupervised instances coherently. In the 2D distillation step, the extracted 3D instance indices help to constrain the correspondence of multi-view images to learn view-consistent semantic embeddings. Consequently, semantics and view-consistent instances are optimized complementarily in a self-bootstrapped manner. However, the 3D semantics in IDSF are still indirectly injected from the 2D plane, which may lead to optimization difficulty incurred by the rendering process of 3DGS. To address this, we further introduce a joint contrastive loss to bridge the 2D-3D discrepancy and boost the learning efficacy.
@@
3D Gaussian Splatting. Recently, 3D scene reconstruction using 3DGS (Kerbl et al. 2023) has gained prominence due to its fast rendering capabilities and impressive visual quality. Unlike the implicit radiance field in NeRF (Mildenhall et al. 2021), 3DGS represents the 3D scene with millions of Gaussians, allowing for more flexible access to the radiance field. Building on the success of 3DGS, several subsequent works have integrated it with 2D foundation models for various applications, such as 3D scene generation, editing, and understanding. Among these methods, one popular trend involves adding constraints from 2D space (e.g., Diffusion models) to enable quality enhancement or promptable editing/generation (Tang et al. 2024b; Yi et al. 2024; Tang et al. 2024a; Yang et al. 2024; Fang et al. 2024). Meanwhile, some concurrent approaches incorporate extra learnable parameters to the original Gaussians for applications beyond 3D reconstruction. For example, some methods (Wu et al. 2024; Zeng et al. 2024; Ren et al. 2023) introduce an extra-temporal dimension into 3DGS for dynamic scene reconstruction, also known as 4D scene synthesis.
@@
Feature Field Distillation. The integration of semantic feature field into radiance field is widely explored in NeRF-based approaches (Zhi et al. 2021; Kobayashi, Matsumoto, and Sitzmann 2022; Fan et al. 2022). Early attempts like Semantic NeRF (Kobayashi, Matsumoto, and Sitzmann 2022) successfully incorporate geometry, appearance, and semantics into the radiance field, showing that precise 3D segmentation can be learned from sparse and noisy 2D annotations. Later, DFF (Kobayashi, Matsumoto, and Sitzmann 2022), LERF (Kerr et al. 2023), and 3DOVS (Liu et al. 2023) distill semantic feature from CLIP (Radford et al. 2021) or DINO (Caron et al. 2021) into NeRF to achieve open-world 3D perception. Recently, these advances from NeRF have been introduced to 3DGS (Lan et al. 2023; Hu et al. 2024; Huang and Yu 2023) to form a more efficient and scalable 3D reconstruction & understanding framework. Beyond CLIP, SAM (Kirillov et al. 2023) is commonly used to generate 2D masks for training views, serving as a weak supervision to improve mask quality. LangSplat (Qin et al. 2023) uses CLIP to supervise the newly-added semantic embeddings and adopt multi-granularity 2D segmentation masks from SAM for fine-grained perception. However, the learned semantic embedding fails to capture view-consistent identity without an effective cross-view constraint. Gaussian Grouping (Ye et al. 2023) assigns each Gaussian a view-consistent ID for cross-view perception. However, this method requires supervision from predefined view-consistent labels, which are generated through complex preprocessing with SAM and tracking models (Cheng et al. 2023). Although most recent methods like CGC (Silva et al. 2024) design contrastive losses to avoid view-consistent 2D masks, the frame-wise SAM masks are still required to group the Gaussians of the same identity. Building upon a 3DGS-based framework, we introduce a bootstrapping 3D clustering strategy to learn view-consistent semantic fields. This algorithm not only eliminates the time-consuming SAM-based 2D mask generation process but also contributes to prompt-unified segmentation in both 2D and 3D scenarios.
@@
In this paper, we propose FreeGS for achieving view-consistent 3D scene understanding without the need of 2D/3D labels. The overall framework of FreeGS is illustrated in Fig. 2. Given only multi-view images as input, the IDentity-coupled Semantic Field (IDSF) is designed to capture both semantic embeddings and view-consistent instance indices. The IDSF is optimized alternately between two stages: Union-space 3D Gaussian Clustering and Multi-level 2D Semantic Distillation. In 3D space, the Union-Space Clustering is used to extract the view-consistent instance indices. Then, the produced indices are projected onto 2D space to regularize the semantic injection with foundational models (Radford et al. 2021; Dong et al. 2023). Additionally, 2D-3D Joint Contrastive Learning is employed to aid in the complementary of 3D geometry and 2D semantics during this bootstrapping process.
@@
where $\mathcal{N}$ denotes the number of Gaussians in the tile and $\alpha^{\prime}_{i}$ represents the Gaussian’s final influence on certain pixel. The transmittance $T_{i}$ is calculated by $T_{i}=\prod_{j=1}^{i-1}(1-\alpha^{\prime}_{j})$. During training, the Gaussians are initialized with the sparse points from SfM (Snavely, Seitz, and Szeliski 2006). Then, they are optimized by constructing a reconstruction loss between ground truth image $I_{GT}$ and the reconstructed image $I$ in various views:
@@
Pixel-level Feature Distillation. The large resolution discrepancy between the rendered feature map (e.g., $1080\times 1440$) and CLIP features (e.g., $14\times 14$) would lead to detail degradation (also known as grid artifacts). We propose pixel-level feature distillation to enhance the details of IDSF. To this end, we exploit an open-vocabulary segmentation model, MaskCLIP (Dong et al. 2023), to extract 2D semantic guidance for pixel-level feature distillation. Specifically, we feed the rendered image $I$ into MaskCLIP and adopt a recent work named FeatUP (Fu et al. 2024) to increase the feature resolution to $224\times 224$, which is represented as $\mathbf{F}_{pix}$. To rectify the resolution misalignment between $\mathbf{F}$ and $\mathbf{F}_{pix}$, we adopt a convolutional layer to downsample $\mathbf{F}$ to the same resolution as $\mathbf{F}_{pix}$, which is denoted as $\hat{\mathbf{F}}$.
@@
Once trained, we can easily extract 3D instance proposals (i.e., clustered groups) and acquire their average features and 3D bounding boxes from the component Gaussians. For each query, we compute relevancy scores between the query embedding and the average feature of each group following LERF (Kerr et al. 2023), and choose the group with the highest score. Then we can generate the corresponding 2D segmentation map or 3D box to achieve novel-view segmentation or 3D object detection. For interactive prompts like click, we are able to directly access the Gaussian according to the click coordinate in 3D space and return the corresponding instance index to produce 2D masks or 3D boxes.
@@
Experiments
Datasets and Metrics
[…]
Implementation Details
We implement the proposed method based on 3DGS  (Kerbl et al. 2023) and conduct all experiments on an Nvidia A800 GPU. We use the cuML library to support all machine learning algorithms in our work, like HDBSCAN and PCA for GPU acceleration. For scene reconstruction, we follow Mini-Splatting (Fang and Wang 2024) for a more uniform spatial distribution of 3D Gaussians. For each scene, the first training phase takes 30k iterations and the second phase for 7k iterations. We set the hyperparameters of $\lambda_{S}$ and $\lambda_{C}$ as 0.1 and 0.05 respectively, and set the feature vector dimension $D=128$. We use the Adam optimizer with a learning rate of 0.0025 and 0.0001 for GS and CNN layers respectively. For $L_{F}$, we set $\gamma=0.3$. For $L_{S}$, we set $K=5$ and $T$ as 0.1% of the total number of Gaussians. We apply $L_{S}$ and $L_{C}$ every 10 iterations for acceleration.
Comparisons with State-of-the-art Methods
Tab. 2 reports the quantitative results on the 3D-OVS dataset. We found that previous methods didn’t split datasets during training, so we retrained the methods with only training views. Note that LERF fails to reconstruct some scenes in this setting, so we use the metric reported in LangSplat. As shown, without any preprocessing, our method achieves comparable overall performance to Gaussian Grouping. Among the five scenes, our method achieves the best performance on “bench” and “sofa”, with a significant improvement of 3.99% and 6.3% in mIoU, respectively.
Open Vocabulary 3D Detection.
Figure 4: Qualitative comparisons of different methods on the ScanNet dataset. We use yellow lines to draw the ground truth 3D bounding boxes. For the two queries in the figure, LangSplat generates bounding boxes as large as the whole scene because of the existence of noise points, so we do not show box visualization for LangSplat. While our method demonstrates much more accurate localization and segmentation.
@@
Table 3: Quantitative comparisons of 3D detection on ScanNet dataset. We report the 3D box’s mIOU, Recall50 and Recall25.
Previous methods fail to encode semantic field and view-consistent identity indices simultaneously in 3DGS or NeRF, hindering their application to 3D perception tasks. To better evaluate the performance on 3D detection, we extend LangSplat to perform 3D detection by generating 3D bounding boxes for Gaussians with high relevancy scores to queries. Tab. 3 illustrates the quantitative comparison results. As shown, without a specific design for the task, our method outperforms LangSplat by a large margin, which further demonstrates the superiority of our method on the 3d scene understanding task. Besides, Fig. 4 shows the visual comparison results. We observe that Langsplat suffers from severe noisy points and cross-view inconsistency, which leads to imprecise box predictions (often as large as the whole scene). In contrast, our clustering-based method builds a more compact and discriminative feature field, successfully producing accurate masks and boxes for each text query.
@@
We conduct ablation studies to analyze the impact of each component in our method. The results are presented in Tab. 4 and Fig. 5. Without clustering, our model can only be queried in 2D space and therefore do not support 3D tasks on ScanNet dataset. Although the performance seems not bad at first sight in this setting, the “Portuguese egg tart” and “mini offroad car” cannot be recognized at all. The 2D-3D joint contrastive learning is primarily designed to stabilize the clustering and pull fragments together. Without $\mathcal{L}_{C}$, the clustering results can sometimes be quite different as training proceeds, and instances are often partitioned into fragments. During experiments on the “figurines” scene, we found an interesting observation that MaskCLIP fails to distinguish between red apple and green apple while CLIP can do that easily, and this is the motivation for the design of $\mathbf{F}_{ins}$. Also, as shown in the figure, $\mathbf{F}_{ins}$ helps our method to recognize “wood”, which indicates that $\mathbf{F}_{ins}$ improves our method’s instance discrimination capability.

### v2

@@ Abstract
Injecting semantics into 3D Gaussian Splatting (3DGS) has recently garnered significant attention. While current approaches typically distill 3D semantic features from 2D foundational models (e.g., CLIP and SAM) to facilitate novel view segmentation and semantic understanding, their heavy reliance on 2D supervision can undermine cross-view semantic consistency and necessitate complex data preparation processes, therefore hindering view-consistent scene understanding. In this work, we present FreeGS, an unsupervised semantic-embedded 3DGS framework that achieves view-consistent 3D scene understanding without the need for 2D labels. Instead of directly learning semantic features, we introduce the IDentity-coupled Semantic Field (IDSF) into 3DGS, which captures both semantic representations and view-consistent instance indices for each Gaussian. We optimize IDSF with a two-step alternating strategy: semantics help to extract coherent instances in 3D space, while the resulting instances regularize the injection of stable semantics from 2D space. Additionally, we adopt a 2D-3D joint contrastive loss to enhance the complementarity between view-consistent 3D geometry and rich semantics during the bootstrapping process, enabling FreeGS to uniformly perform tasks such as novel-view semantic segmentation, object selection, and 3D object detection. Extensive experiments on LERF-Mask, 3D-OVS, and ScanNet datasets demonstrate that FreeGS performs comparably to state-of-the-art methods while avoiding the complex data preprocessing workload.
Codehttps://github.com/wb014/FreeGS
@@ Introduction
Figure 1: We present a novel framework, FreeGS, to inject 2D semantics (Refer to caption) into 3DGS (Refer to caption), without the need of any 2D labels. After a 2D-3D collaborative bootstrapping learning strategy, the model can support versatile applications, such as novel-view 2D segmentation, open-vocabulary 3D detection, and interactive object selection.
@@
To resolve these challenges, we present FreeGS, an unsupervised 3DGS framework for view-consistent 3D scene understanding that uses only raw images as input, without requiring any labels. The core of FreeGS is the IDentity-coupled Semantic Field (IDSF), which holds both semantic embeddings and view-consistent instance indices for each Gaussian. With only multi-view images as input, we optimize IDSF through a bootstrapping process that alternates between the Union-space 3D Gaussian Clustering phase and the Multi-level 2D Semantic Distillation phase. In the 3D clustering step, semantic information collaborates with the intrinsic properties of Gaussians (e.g., geometry and appearance) to extract unsupervised instances coherently. In the 2D distillation step, the extracted 3D instance indices help to constrain the correspondence of multi-view images to learn view-consistent semantic embeddings. Consequently, semantics and view-consistent instances are optimized complementarily in a self-bootstrapped manner. However, the 3D semantics in IDSF are still indirectly injected from the 2D plane, which may lead to optimization difficulty incurred by the rendering process of 3DGS. To address this, we further introduce a joint contrastive loss to bridge the 2D-3D discrepancy and boost the learning efficacy.
@@
3D Gaussian Splatting. Recently, 3D scene reconstruction using 3DGS (Kerbl et al. 2023) has gained prominence due to its fast rendering capabilities and impressive visual quality. Unlike the implicit radiance field in NeRF (Mildenhall et al. 2021), 3DGS represents the 3D scene with millions of Gaussians, allowing for more flexible access to the radiance field. Building on the success of 3DGS, several subsequent works have integrated it with 2D foundation models for various applications, such as 3D scene generation, editing, and understanding. Among these methods, one popular trend involves adding constraints from 2D space (e.g., Diffusion models) to enable quality enhancement or promptable editing/generation (Tang et al. 2024b; Yi et al. 2024; Tang et al. 2024a; Yang et al. 2024; Fang et al. 2024). Meanwhile, some concurrent approaches incorporate extra learnable parameters to the original Gaussians for applications beyond 3D reconstruction. For example, some methods (Wu et al. 2024b; Zeng et al. 2024; Ren et al. 2023) introduce an extra-temporal dimension into 3DGS for dynamic scene reconstruction, also known as 4D scene synthesis.
@@
Feature Field Distillation. The integration of semantic feature field into radiance field is widely explored in NeRF-based approaches (Zhi et al. 2021; Kobayashi, Matsumoto, and Sitzmann 2022; Fan et al. 2022). Early attempts like Semantic NeRF (Kobayashi, Matsumoto, and Sitzmann 2022) successfully incorporate geometry, appearance, and semantics into the radiance field, showing that precise 3D segmentation can be learned from sparse and noisy 2D annotations. Later, DFF (Kobayashi, Matsumoto, and Sitzmann 2022), LERF (Kerr et al. 2023), and 3DOVS (Liu et al. 2023) distill semantic feature from CLIP (Radford et al. 2021) or DINO (Caron et al. 2021) into NeRF to achieve open-world 3D perception. Recently, these advances from NeRF have been introduced to 3DGS (Lan et al. 2023; Hu et al. 2024; Huang and Yu 2023) to form a more efficient and scalable 3D reconstruction & understanding framework. Beyond CLIP, the Segment Anything Model (SAM) (Kirillov et al. 2023) is commonly used to generate 2D masks for training views, serving as a weak supervision to improve mask quality. LangSplat (Qin et al. 2023) uses CLIP to supervise the newly-added semantic embeddings and adopt multi-granularity 2D segmentation masks from SAM for fine-grained perception. However, the learned semantic embedding fails to capture view-consistent identity without an effective cross-view constraint. Gaussian Grouping (Ye et al. 2023) assigns each Gaussian a view-consistent ID for cross-view perception. However, this method requires supervision from predefined view-consistent labels, which are generated through complex preprocessing with SAM and tracking models (Cheng et al. 2023). Some concurrent works also attempt to enhance the feature field with various motivations. For example, N2F2 (Bhalgat et al. 2024) improves the feature field to be scale-aware and fairly reduces the consuming time of training and testing. CGC (Silva et al. 2024) and OpenGaussian (Wu et al. 2024a) both design contrastive losses to avoid the need for view-consistent 2D labels. However, all of them still need frame-wise SAM masks to group the Gaussians of the same identity. Building upon a 3DGS-based framework, we introduce a bootstrapping 3D clustering strategy to learn view-consistent semantic fields. This algorithm not only eliminates the time-consuming SAM-based 2D mask generation process but also contributes to prompt-unified segmentation in both 2D and 3D scenarios.
@@
In this paper, we propose FreeGS for achieving view-consistent 3D scene understanding without the need for 2D/3D labels. The overall framework of FreeGS is illustrated in Fig. 2. Given only multi-view images as input, the IDentity-coupled Semantic Field (IDSF) is designed to capture both semantic embeddings and view-consistent instance indices. The IDSF is optimized alternately between two stages: Union-space 3D Gaussian Clustering and Multi-level 2D Semantic Distillation. In 3D space, the Union-Space Clustering is used to extract the view-consistent instance indices. Then, the produced indices are projected onto 2D space to regularize the semantic injection with foundational models (Radford et al. 2021; Dong et al. 2023). Additionally, 2D-3D Joint Contrastive Learning is employed to aid in the complementary of 3D geometry and 2D semantics during this bootstrapping process.
@@
where $\mathcal{N}$ denotes the number of Gaussians in the tile and $\alpha^{\prime}_{i}$ represents the Gaussian’s final influence on certain pixels. The transmittance $T_{i}$ is calculated by $T_{i}=\prod_{j=1}^{i-1}(1-\alpha^{\prime}_{j})$. During training, the Gaussians are initialized with the sparse points from SfM (Snavely, Seitz, and Szeliski 2006). Then, they are optimized by constructing a reconstruction loss between ground truth image $I_{GT}$ and the reconstructed image $I$ in various views:
@@
Here, $\lambda=0.2$ and the pixels of reconstructed image $I$ are obtained by Eq. 1. The explicit 3D representations with Gaussians and the tile-based rasterizer collaborate to improve the rendering efficiency and enable flexible scene editing via directly manipulating the desired Gaussians.
@@
Pixel-level Feature Distillation. The large resolution discrepancy between the rendered feature map (e.g., $1080\times 1440$) and CLIP features (e.g., $14\times 14$) would lead to detail degradation (also known as grid artifacts). Therefore, we propose pixel-level feature distillation to enhance the details of IDSF. To this end, we exploit MaskCLIP (Dong et al. 2023) as a feature extractor to extract 2D spatially vision-text-aligned features from the origin CLIP visual encoder, for pixel-level feature distillation. Specifically, we feed the rendered image $I$ into MaskCLIP and adopt a recent feature super-resolution work named FeatUP (Fu et al. 2024) to increase the feature resolution to $224\times 224$, which is represented as $\mathbf{F}_{pix}$. To rectify the resolution misalignment between $\mathbf{F}$ and $\mathbf{F}_{pix}$, we also adopt a convolutional layer to downsample $\mathbf{F}$ to the same resolution as $\mathbf{F}_{pix}$, which is denoted as $\hat{\mathbf{F}}$.
@@
Specifically, we abandon the previous process of querying separately under each view, and instead propose a match-by-propose approach as an advanced replacement. Once trained, we can easily extract 3D instance proposals (i.e., clustered groups) and acquire their average features and 3D bounding boxes from the component Gaussians. For each query, we compute relevancy scores between the query embedding and the average feature of each group following LERF (Kerr et al. 2023), and choose the group with the highest score. Then, based on the acquired 3D proposals, we can easily generate the corresponding 2D segmentation map for novel-view segmentation and 3D box for 3D object detection. For interactive prompts like click, we are able to directly access the Gaussian according to the click coordinate in 3D space and return the corresponding instance index to produce 2D masks or 3D boxes.
Experiments
Datasets and Metrics
Datasets. To evaluate the effectiveness of our method, we conduct experiments on two widely used datasets: LERF-Mask (Ye et al. 2023) and 3D-OVS (Liu et al. 2023). LERF-Mask dataset consists of three scenes with dense annotations. These scenes are categorized as “long-tail scenes” of LERF-Localization dataset (Kerr et al. 2023), which presents great challenges for scene understanding tasks, especially in unsupervised settings. 3D-OVS dataset contains uncommon objects like “Gundam” and “Pikachu”, posing even greater challenges to the task. We select five scenes following the criteria of  (Qin et al. 2023). Except for novel-view perception, we evaluate the proposed method on 3D object detection and establish a benchmark on ScanNet (Dai et al. 2017), which is a widely used dataset with complex indoor scenes for 3D scene understanding. In detail, we conduct experiments on three scenes in ScanNet: scene0060, scene0381, and scene0462. For each scene, we sample 10% views for reconstruction, with the sampling strategy based on image sharpness to avoid motion blur.
@@
Figure 4: Qualitative comparisons of different methods on “Scene0462” of the ScanNet dataset. We use yellow lines to draw the ground truth 3D bounding boxes. For the two queries in the figure, LangSplat generates bounding boxes as large as the whole scene because of the existence of noise points, so we do not show box visualization for LangSplat. While our method demonstrates much more accurate localization and segmentation.
@@
Table 3: Quantitative comparisons of open vocabulary 3D detection on the ScanNet dataset. We report the 3D box’s mIOU, Recall50 and Recall25.
Implementation Details
We implement the proposed method based on 3DGS  (Kerbl et al. 2023) and conduct all experiments on an Nvidia A800 GPU. We use the cuML library (Raschka, Patterson, and Nolet 2020) to support all machine learning algorithms in our work, such as HDBSCAN and PCA, for GPU acceleration. For scene reconstruction, we follow Mini-Splatting (Fang and Wang 2024) for a more uniform spatial distribution of 3D Gaussians. For each scene, the first training phase takes 30k iterations and the second phase for 7k iterations. We set the hyperparameters of $\lambda_{S}$ and $\lambda_{C}$ as 0.1 and 0.05 respectively, and set the feature vector dimension $D=128$. We use the Adam optimizer with a learning rate of 0.0025 and 0.0001 for GS and CNN layers respectively. For $L_{F}$, we set $\gamma=0.3$. For $L_{S}$, we set $K=5$ and $T$ as 0.1% of the total number of Gaussians. We apply $L_{S}$ and $L_{C}$ every 10 iterations for acceleration. For hyperparameters used in HDBSCAN, we set $min\_samples=20$ in experiments on LERF-Mask, and $min\_samples=60$ in experiments on 3D-OVS and ScanNet.
Comparisons with State-of-the-art Methods
Tab. 2 reports the quantitative results on the 3D-OVS dataset. We found that previous methods did not split datasets during training, so we retrained the methods with only training views and tested them on testing views to evaluate generalization. Note that LERF fails to reconstruct some scenes in this setting, so we use the metric reported in LangSplat. As shown, without any preprocessing, our method achieves comparable overall performance to Gaussian Grouping. Among the five scenes, our method achieves the best performance on “bench” and “sofa” scenes, with a significant improvement of 3.99% and 6.3% in mIoU metric, respectively.
  Method Preprocess(m) Training(m) Total(m)
LERF 0 54 54
G-Group 24 68 92
LangSplat 141 105 246
Ours 0 62 62
Table 4: Efficiency comparisons of different methods on the “ramen” scene of the LERF-Mask dataset. Our method significantly surpasses previous SAM-based approaches in efficiency.
Open Vocabulary 3D Detection. Previous methods fail to encode semantic field and view-consistent identity indices simultaneously in 3DGS or NeRF, hindering their application to 3D perception tasks. To better evaluate the performance on the 3D detection task, we extend LangSplat to perform 3D detection by generating 3D bounding boxes for Gaussians with high relevancy scores to text queries. Tab. 3 illustrates the quantitative comparison results. As shown, without a specific design for the task, our method outperforms LangSplat by a large margin, which further demonstrates the superiority of our method on the 3d scene understanding task. Besides, Fig. 4 shows the visual comparison results. We observe that Langsplat suffers from severe noisy points and cross-view inconsistency, which leads to imprecise box predictions (often as large as the whole scene). In contrast, our clustering-based method builds a more compact and discriminative feature field, successfully producing accurate masks and boxes for each text query.
Preprocessing & Training Time. Tab. 4 illustrates the efficiency comparisons on the “ramen” scene of the LERF-Mask dataset. Although LERF avoids the preprocessing phase and achieves the fastest, its accuracy is notably low as detailed in Tab. 1. For SAM-based methods G-Group and LangSplat, preprocessing times are 24 and 141 minutes respectively, accounting for over 35% and 134% of their training time. The complex and time-consuming preprocessing significantly impairs the models’ efficiency and practical applicability.
@@
We conduct comprehensive ablation studies to analyze the impact of each component in our method. The results are presented in Tab. 5 and Fig. 5. Without clustering, our model can only be queried in 2D space and therefore do not support 3D tasks on the ScanNet dataset. Although the performance seems not bad at first sight in this setting, the “Portuguese egg tart” and “mini offroad car” cannot be recognized at all. The 2D-3D joint contrastive learning is primarily designed to stabilize the clustering and pull instance fragments together. Without $\mathcal{L}_{C}$, the clustering results can sometimes be quite different as training proceeds, and instances are often partitioned into fragments. During experiments on the “figurines” scene, we found an interesting observation that MaskCLIP fails to distinguish between “red apple” and “green apple” while CLIP can do that successfully, and this is in fact our motivation for the design of $\mathbf{F}_{ins}$. Also, as shown in the figure, $\mathbf{F}_{ins}$ helps our method to recognize “wood”, which indicates that $\mathbf{F}_{ins}$ improves our method’s instance discrimination capability.
@@
Figure 6: Visualization of interactive object selection results with point or box as prompts. Scenes are “bed” and “room” of the 3D-OVS dataset.